OpenStudy (mrhoola):

Hwo do you convert the product to polar form : P = (12 +j5)(1<30 degrees)

OpenStudy (mrhoola):

Note 1<30 degrees is a phasor.. AND you CANNOT use a calculator also its easy if you have a calculator but when you convert the rectangular part to its phase angle inverse tan = (5/12) , I dont know how the heck am I suppose to know the value in degrees . conversely converting the phasor to rectangular is difficult with out a calculator

OpenStudy (freckles):

phaser? You can give me example of what that means exactly?

OpenStudy (mrhoola):

phasor is equivalent to polar coordinates

OpenStudy (freckles):

And I'm assuming your j is my i that is, j is the actual imaginary unit

OpenStudy (mrhoola):

the work is obvious but NOT clean . thats the problem becuase we cannot use calculators on the exam , just want to clarify that

OpenStudy (mrhoola):

yes

OpenStudy (freckles):

so is that second factor just: \[\cos(30)+j \sin(30)\]

OpenStudy (freckles):

or in the other form: \[e^{j \frac{\pi}{6}}\]

OpenStudy (mrhoola):

(12 +j5)*(sqrt(3)/2 +j(1/2)) .. applying multiplication makes it a hair ball ,

OpenStudy (freckles):

\[P=(12+j5)e^{\frac{\pi}{6}j}\] maybe we can also convert the other one to polar form then find the product

OpenStudy (mrhoola):

ok but , inversse tan = (12/5)

OpenStudy (mrhoola):

how do you implement that

OpenStudy (freckles):

you cannot just leave theta as arctan(5/12) also it is y/x inside tan

OpenStudy (freckles):

?

OpenStudy (mrhoola):

oh yeah my bad

OpenStudy (freckles):

\[P=\sqrt{12^2+5^2}e^{\arctan(\frac{5}{12})j} \cdot e^{\frac{\pi}{6} j} \\ P=\sqrt{169} e^{\arctan(\frac{5}{12})j}e^{\frac{\pi}{6} j} \\ P=13 e^{(\arctan(\frac{5}{12})+\frac{\pi}{6})j}\]

OpenStudy (freckles):

would that not work?

OpenStudy (freckles):

and just to be clear that factor rewrite was correct right? 1<30 degrees means e^(pi/6 i) ?

OpenStudy (mrhoola):

I got another method hold on let me write it . i thhink it workds

OpenStudy (freckles):

oops j not i

OpenStudy (mrhoola):

(12+ j5)(cos(30) + sin(30)) = 12cos(30) -sin()30 +j (5cos(30)+12sin(30)).....can you help me simplify that

OpenStudy (mrhoola):

maybe if you conver cos to sin ..... sin (theta) = cos(theta - 90 ) ??? is that valid

OpenStudy (mrhoola):

no no its not never mind

OpenStudy (freckles):

\[(12+5j)(\cos(\frac{\pi}{6})+j \sin(\frac{\pi}{6})) \\ =12(\cos(\frac{\pi}{6})+j \sin(\frac{\pi}{6}))+5j(\cos(\frac{\pi}{6})+j \sin(\frac{\pi}{6})) \\ =12 \cos(\frac{\pi}{6})+12 j \sin(\frac{\pi}{6})+5j \cos(\frac{\pi}{6})+5j^2 \sin(\frac{\pi}{6}) \\ \text{ and } j^2=-1 \\ \text{ so we have } \\ =12 \cos(\frac{\pi}{6})+12 j \sin(\frac{\pi}{6})+5j \cos(\frac{\pi}{6})-5 \sin(\frac{\pi}{6}) \\\] hmmm cos(x)=sin(pi/2-x) so cos(pi/6)=sin(pi/2-pi/6)=sin(2pi/6)=sin(pi/3) and sin(pi/6)=cos(pi/2-pi/6)=cos(pi/3) hmm... So we have: \[12e^{ \frac{\pi}{6} j} +5j \sin(\frac{\pi}{3})-5 \cos(\frac{\pi}{3})\] is that what you wanted to do...

OpenStudy (mrhoola):

yes its close enough, you still have it in rectangular form - it still needs to be converted to polar . from what you have shown me - I cant imagine any other way to do this problem .

OpenStudy (freckles):

\[12e^{ \frac{\pi}{6} j} -5 \cos(\frac{\pi}{3})+5 j \sin(\frac{\pi}{3}) \\ 12e^{\frac{\pi}{6} j}-5(\cos(\frac{\pi}{3})-j \sin(\frac{\pi}{3}) ) \\ 12e^{\frac{\pi}{6} j } -5 (\cos(\frac{\pi}{3})+j \sin(-\frac{\pi}{3})) \\ 12e^{\frac{\pi}{6}j} -5(\cos(\frac{-\pi}{3})+j \sin(-\frac{\pi}{3}))\] since cos(-x)=cos(x) and sin(-x)=-sin(x)

OpenStudy (freckles):

now that one part should be ready to go

OpenStudy (mrhoola):

Thank you , I have to double check with my professor . I will get back to you eventually

OpenStudy (freckles):

but you do see that one thing following the difference sign can be written in polar form right?

OpenStudy (freckles):

\[12e^{\frac{\pi}{6} j}-5e^{\frac{-\pi}{3} j}\]

OpenStudy (freckles):

can be written as a difference in polar form

OpenStudy (mrhoola):

To be honest I have never seen E raised to x/ a .. That is completely new to me

OpenStudy (freckles):

honestly though I kinda would have stuck with the above answer \[P=\sqrt{12^2+5^2}e^{\arctan(\frac{5}{12})j} \cdot e^{\frac{\pi}{6} j} \\ P=\sqrt{169} e^{\arctan(\frac{5}{12})j}e^{\frac{\pi}{6} j} \\ P=13 e^{(\arctan(\frac{5}{12})+\frac{\pi}{6})j}\] But you are totally right I don't know what goes in your course

OpenStudy (freckles):

oh what do you call polar form \[e^{ i \theta} \text{ is the same as } \cos(\theta)+i \sin(\theta) \]

OpenStudy (freckles):

I call exp(i theta) polar form

OpenStudy (freckles):

http://web.eecs.umich.edu/~aey/eecs206/lectures/phasor.pdf these use that imaginary unit j so I assume this is the engineering version of the trig class I took

OpenStudy (freckles):

They call polar form the same thing I do

OpenStudy (mrhoola):

Thank you for the resourceful document - I will study this !!